write an equation of the line that passes through the given points...
(-2,8) and (1,2)

Answer by Alan3354(69443)

write an equation of the line that passes through the given points...
(-2,8) and (1,2)
-------
Here's 1 way to do that:
Find the slope, m
m = diffy/diffx = -6/3 = -2
---
y-y1 = m*(x-x1) where (x1,y1) is either point
y-2 = -2*(x-1)
y = -2x + 4
======================
Here's another way:

| x   y   1|
|-2   8   1| = 0
| 1   2   1|


x*(8-2) - y*(-2-1) + 1*(-4-8) = 0
6x + 3y = 12
2x + y = 4